• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name jpchen

    Log entry time 17:53:55 on July 2,2010

    Entry number 329474

    keyword=Target cooldown completed

    Target cool-down completed.
    Coolant return temperature (loop 2) is 27 K and will continue to
    decrease slowly. Supply is 24K (Hall A side) and 17.6 K (ESR side).
    They are also slowly decreasing.
    It will take another hour or so to stabalize.
    Scattering chamber vacuum only dropped to ~1.5*10-5. May decrease a little more with continued cooling. Not ideal (usually can reach 10-6). Hopefully it will stay low.
    I set alarm for vacuum to be HI=5*10-5, HIHI=1*10-4, for return
    temperature HI=30, HIHI=32.


    FIGURE 1

    FIGURE 2